Consider a 0.400 M CH3COOH solution (Ka=1.8*10^-5) and determine the [H3O+] concentration at equilibrium
Answer #1

We know

Ka = [CH3COO-][H3O+]/[CH3COOH]

Let X= [CH3COO-][H3O+]

1.8×10^-5= X * X/ (0.4-X)

X2 = 1.8×10^-5 (0.4-x)

On solving this quadratic equation we know that get

X=[H3O+] = 0.0002674 = 2.674×10^-4